OpenStudy (anonymous):

/4x+8/=9 is the equation always,sometimes,never true

OpenStudy (the_fizicx99):

" /4x + 8/ " why is there division bars? Are you missing something?

OpenStudy (anonymous):

know there sopose to be bars in front

OpenStudy (the_fizicx99):

You mean |4x + 8| = 9 ?

OpenStudy (anonymous):

yes

OpenStudy (the_fizicx99):

Absolute values so to drop them you make another equation, 4x + 8 = 9 and 4x + 8 = -9 solve

OpenStudy (the_fizicx99):

The opposite sign, its positive so it becomes negative

OpenStudy (anonymous):

is it always

OpenStudy (the_fizicx99):

Hm, well its absolute value so it should be 9 and -9